A stone is dropped into a pool of water. The radius of the circular ripple formed increases at 4 π‘š/𝑠. Calculate the rate at which the area of the ripple is increasing when the radius is 6m.


1
Expert's answer
2021-11-21T17:33:24-0500

"\frac{dr}{dt}=4m/s"


"A=πr²=>\frac{dA}{dr}=2πr"


"\frac{dA}{dt}=\frac{dA}{dr}•\frac{dr}{dt}"


"=>\frac{dA}{dt}=2πr•4=8πrm²/s"


when "r=6m"


"\frac{dA}{dt}=8π(6)=48πm²/s"